a box contains nickels, dimes and pennies worth $5.56. The number of pennies is one more than twice the number of dimes and there is an equal number of pennies and nickels. how many nickels are there in the box?
Answer by ankor@dixie-net.com(22740)

let n = no. of nickels
let d = no. of dimes
let p = no. of pennies
:
Write an equation for each statement
:
a box contains nickels, dimes and pennies worth $5.56.
.01p + .05n + .10d = 5.56
:
The number of pennies is one more than twice the number of dimes
p = 2d+1
:
and there is an equal number of pennies and nickels.
p = n
therefore
n = 2d+1
In the first equation replace p and n with (2d+1)
.01(2d+1) + .05(2d+1) + .10d = 5.56
.02d + .01 + .10d + .05 + .10d = 5.56
.22d + .06 = 5.56
.22d = 5.50
d = 5.5/.22
d = 25 dimes
;
How many nickels are there in the box?
n = 2(25) + 1
n = 51 dimes
:
:
See if this adds up
.01(51) + .05(51) + .10(25) =
.51 + 2.55 + 2.50 = 5.56